OpenStudy (anonymous):

Prove that if you keep adding the digits of any number and get the answer 3,6 or 9 then the number is divisible by 3. I used to know this proof but my brain circuits cant reach it anymore!

OpenStudy (amistre64):

the rule of 9 right?

OpenStudy (amistre64):

15 = 6 which divides by 3 i know that the factors of 9 add up to 9; might be a correlary to that :)

OpenStudy (amistre64):

9*11 = 99 tho so it dont count :)

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so get an odd number you have to add by at least one odd right maybe?

OpenStudy (amistre64):

but 6 is even lol

OpenStudy (amistre64):

i cant recall if i ever knew this :)

OpenStudy (anonymous):

i did but its completely gone

OpenStudy (anonymous):

if u take a 3 digit number abc then if we can prove that a + b + c = n/3 where n is an positive integer that would prove it true for 3 dgts

OpenStudy (anonymous):

a number N can be written a0 + 10a1 +10^2a2 etc. 10 = 1 mod 9 so 10^r = 1 mod 9 so N = a0 + a1 + a2... mod9 and relsult follows (ie divisible by 3 iff sum of digits is).

OpenStudy (anonymous):

Jimmy I have a question I need to find the domain of a rational function f(x)=3x-8 __________ 4x+20